### v1

# Dynamic Pre-training: Towards Efficient and Scalable All-in-One Image Restoration

Abstract
All-in-one image restoration tackles different types of degradations with a unified model instead of having task-specific, non-generic models for each degradation. The requirement to tackle multiple degradations using the same model can lead to high-complexity designs with fixed configuration that lack the adaptability to more efficient alternatives. We propose DyNet, a dynamic family of networks designed in an encoder-decoder style for all-in-one image restoration tasks. Our DyNet can seamlessly switch between its bulkier and lightweight variants, thereby offering flexibility for efficient model deployment with a single round of training. This seamless switching is enabled by our weights-sharing mechanism, forming the core of our architecture and facilitating the reuse of initialized module weights. Further, to establish robust weights initialization, we introduce a dynamic pre-training strategy that trains variants of the proposed DyNet concurrently, thereby achieving a 50% reduction in GPU hours. To tackle the unavailability of large-scale dataset required in pre-training, we curate a high-quality, high-resolution image dataset named Million-IRD having 2M image samples. We validate our DyNet for image denoising, deraining, and dehazing in all-in-one setting, achieving state-of-the-art results with 31.34% reduction in GFlops and a 56.75% reduction in parameters compared to baseline models. The source codes and trained models are available at https://github.com/akshaydudhane16/DyNet.
Keywords:
All-in-one Image Restoration Foundation model
1 Introduction
The image restoration (IR) task seeks to improve low-quality input images. Despite several advancements in IR, diverse degradation types and severity levels present in images continue to pose a significant challenge. The majority of existing methods [46, 31, 13, 8, 15, 52, 3, 14, 47] learn image priors implicitly, requiring separate network training for different degradation types, levels, and datasets. Further, these methods require a prior knowledge of image degradation for effective model selection during testing, thereby lacking generality to cater diverse degradations.
All-in-one restoration aims to restore images with an unknown degradation using a single unified model. Recent advancements, such as AirNet [26] and PromptIR [36], have addressed the all-in-one restoration challenge by employing contrastive learning and implicit visual prompting techniques, respectively. Specifically, the state-of-the-art PromptIR [36] employs an implicit prompting to learn degradation-aware prompts on the decoder side, aiming to refine the decoder features. This approach, while intriguing, does not extend to the refinement of encoder features, which are left unprocessed. Despite the interesting idea of implicit prompting, it poses a considerable challenge for deployment due to its poor computational efficiency with 37M parameters and 243 GFlops to process a single 224$\times$224 sized image. The practical application of such models is therefore challenging due to the high computational requirements, especially in resource-constraint hand-held devices. However, choosing lightweight models invariably leads to a trade-off between accuracy and efficiency [56, 28]. Most of these models aim to reduce parameter counts by strided convolutions and feature channel splitting [23, 28, 22]. Some strategies also operate within the frequency domain to reduce the computational demands linked to attention mechanisms [62] or non-local operations [18] while few methods like [9, 49] partition the feature space for efficient processing, or split the attention across dimensions to improve the computational efficiency [61, 9].
Towards optimizing all-in-one IR efficiency without sacrificing performance, in this paper, we introduce a novel weights-sharing mechanism. In this scheme, the weights of a network module are shared with its subsequent modules in a series. This approach substantially reduces the number of parameters, resulting in a more streamlined network architecture. We deploy the proposed weights-sharing mechanism in an encoder-decoder style architecture named Dynamic Network (DyNet). Apart from the computational efficiency, the proposed DyNet provides exceptional flexibility as by merely changing the module weights reuse frequency, one can easily adjust the network depth and correspondingly switch between its bulkier and light-weight variants during training.
While DyNet offers great adaptability with orders of speedup, a common challenge with lightweight networks is the potential compromise on overall accuracy. To counteract this, we show that a large-scale pre-training strategy can effectively improve the performance of our lightweight models. By initializing the network with weights derived from the pre-training, the model benefits from a strong foundation, which can lead to enhanced performance even with a reduced parameter count. However, large-scale pre-training is computationally intensive and requires significant GPU hours.
Figure 1: Left: At any given encoder-decoder level, (a) Series of transformer blocks in existing PromptIR [36], (b) The proposed transformer block weights sharing mechanism in our DyNet-L. We initialize one transformer block at each encoder-decoder level and shared its weights with the subsequent $N$ transformer blocks. (c) Our lightweight variant DyNet-S. Here, the number of transformer blocks are halved and weights of the first transformer block are shared with $N/2$ subsequent transformer blocks. Right: A plot of average PSNR in All-in-one restoration setting vs GFlops and parameters (in millions). The proposed DyNet-S achieves boost of 0.59 dB with 31.34% reduction in GFlops and a 56.75% reduction in parameters compared to the baseline PromptIR [36].
Therefore, we introduce an efficient and potent dynamic pre-training strategy capable of training both bulkier and lightweight network variants within a single pre-training session, thereby achieving significant savings of 50% in GPU hours. Complementing this, we have compiled a comprehensive, high-quality, high-resolution dataset termed Million-IRD, consisting of 2 million image samples. Our main contributions are as follows:
  • We propose DyNet, a dynamic family of networks for all-in-one image restoration tasks. DyNet offers easy switching between its bulkier and light-weight variants. This flexibility is made possible through a weight-sharing strategy, which is the central idea of our network design, allowing for the efficient reuse of initialized module weights.
  • We introduce a Dynamic Pre-training strategy, a new approach that allows the concurrent large-scale pre-training of both bulkier and light-weight network variants within a single session. This innovative strategy reduces GPU hours significantly by 50%, addressing the pressing challenge of resource-intensive model pre-training on a large-scale.
  • We curate a comprehensive pre-training dataset comprising 2 million high-quality, high-resolution, meticulously filtered images. From this dataset, we extract 8 million non-overlapping high-resolution patches, each sized 512$\times$512, utilized for the pre-training of variants of the proposed DyNet.
Through the synergy of these techniques, our proposed DyNet achieves an average gain of 0.82 dB for image de-noising, deraining, and dehazing within an all-in-one setting, with 31.34% reduction in GFlops and a 56.75% reduction in network parameters compared to the baseline; refer Fig. 1.
2 Related Work
Image restoration seeks to restore images from their degraded low-quality versions, with the restoration process varying considerably across different tasks. Research in image restoration has predominantly concentrated on addressing single degradation challenges [52, 40, 11, 53, 38, 58, 45, 34, 60]. Although significant progress has been made in single degradation restoration techniques, research in multi-degradation restoration using a unified model is still relatively under-explored. Multi-degradation restoration, however, provides more practical advantages as the information about input degradation type is not required to select the task-specific model and it also enhances model storage efficiency over single-degradation restoration.
The critical challenge in multi-task image restoration is developing a single model capable of addressing various types of degradation and precisely restoring low-quality images. The IPT model [7], for instance, relies on prior knowledge of the corruption affecting the input image, utilizing a pre-trained transformer backbone equipped with distinct encoders and decoder heads tailored for multi-task restoration. However, in blind image restoration, we do not have such prior information about the degradation. In this direction, Li et al. [26] propose a unified network for multiple tasks such as denoising, deraining, and dehazing, employing an image encoder refined through contrastive learning. However, it requires a two-stage training process, where the outcome of contrastive learning relies on selecting positive and negative pairs and the availability of large data samples. Further, motivated by easy implementation and broad applicability of prompts, PromptIR [36] introduces a concept where degradation-aware prompts are learned implicitly within an encoder-decoder architecture for all-in-one image restoration. PromptIR applies implicit prompting to the decoder side to refine the decoder features. However, it does not extend to the refinement of encoder features, which are left unprocessed.
Although the above-discussed methods tackle multi-task image restoration, they suffer from low efficiency due to a large number of parameters and substantial GFlops consumption. Consequently, deploying these models in environments where efficiency is crucial factor becomes difficult due to their demanding computational needs. However, opting for lightweight models introduces an inevitable accuracy-efficiency trade-off [56]. The primary cause is that many current methods focus on reducing parameter counts through techniques like strided convolutions and feature channel splitting [23, 28, 22]. Additionally, few approaches work within the frequency domain to reduce the computational burden associated with attention mechanisms [62] or non-local operations [18]. Furthermore, methods like [9, 49] partition the feature space for efficient processing, while [61, 9] split the attention across dimensions to improve the computational efficiency. In contrast, our proposed approach focuses on an efficient and scalable all-in-one image restoration model that incorporates a module weights-sharing mechanism. We demonstrate that strategic fundamental adjustments to the network architecture result in substantial performance improvements.
Figure 2: The overall pipeline of the proposed Dynamic Network (DyNet) for all-in-one image restoration. DyNet enhances a low-quality input image using a four-level encoder-decoder architecture. A distinctive aspect of DyNet lies in its weight-sharing strategy: at each level, the initial transformer block’s weights are sequentially passed to subsequent blocks, significantly reducing the network’s parameters and enhancing its flexibility. This approach allows for easy modification of DyNet’s complexity, toggling between more bulkier or lightweight variants by adjusting the frequency of weight sharing among the blocks at each encoder-decoder level. Moreover, we maintain the encoder-decoder feature consistency by implicitly learning degradation-aware prompts at skip connections rather than on the decoder side as in PromptIR [36].
3 Proposed Approach
Existing state-of-the-art all-in-one image restoration (IR) approaches incur a high computational footprint and do not provide the flexibility to support models with variable depths during training. Towards more accurate, but lightweight and flexible architecture design for all-in-one IR, this work proposes a Dynamic Network Architecture (Sec. 3.1) and a Dynamic Pretraining Strategy (Sec. 3.2).
3.1 Dynamic Network (DyNet) Architecture
Our lightweight architecture is based on a novel weight-sharing mechanism for IR models. This mechanism allows the network module’s weights to be reused across subsequent modules in sequence, thereby significantly reducing the total number of parameters and leading to a more efficient network structure. We implement this weight-sharing approach within an encoder-decoder style architecture, which we term the Dynamic Network (DyNet). Within DyNet, at each encoder-decoder level, module weights are shared across a pre-specified number of subsequent modules. This not only enhances computational efficiency but also grants remarkable flexibility to the architecture. By simply altering the frequency of weight sharing, users can easily modify the network’s depth, seamlessly transitioning between bulkier or lightweight variants of DyNet.
3.1.1 Overall Pipeline.
The overall pipeline of the proposed DyNet is shown in Fig. 2. The proposed DyNet begins by extracting low-level features $\mathbf{F_{0}}$in $\mathbb{R}^{H\times W\times C}$ from a given degraded input image $\mathbf{{I}}$in $\mathbb{R}^{H\times W\times 3}$ through a convolution operation. Subsequently, the feature embedding undergoes a 4-level hierarchical encoder-decoder, gradually transforming into deep latent features $\mathbf{F}_{l}$in $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$. We utilize the existing Restormer’s [52] Transformer block as a fundamental feature extraction module. For every level of the encoder and decoder, we initialize the weights for a first transformer block, which are subsequently reused across the subsequent blocks in a series up to the specified frequency at that level. For example, at a given encoder/decoder level, $w^{1}$ denotes weights of the first transformer block, we can define the output $\mathbf{F}_{out}$ of that encoder/decoder level, given input features $\mathbf{F}_{in}$ as,
$\mathbf{F}_{out}$ $=w^{1}(\mathbf{F}_{in})\qquad\qquad for\,\,b=1\qquad\qquad$
[…]
3exOcFFL
b=2:f
[…]
forb=2
[…]
3exOcFFL
[…]
absent
[…]
3exOcFFL
[…]
We gradually increase the reuse frequency of the transformer block from the top level to the bottom level, thereby increasing the network depth. The multi-level encoder systematically reduces spatial resolution while increasing channel capacity, facilitating the progressive extraction of low-resolution latent features from a high-resolution input image. Subsequently, the multi-level decoder gradually restores the high-resolution output from these low-resolution latent features. To enhance the decoding process, we integrate prompt blocks [36], which learn degradation-aware prompts implicitly through the prompt generation module (PGM) followed by the prompt-interaction module (PIM). Unlike the existing PromptIR [36], our approach incorporates degradation-aware implicit prompt blocks at skip connections. This implicit prompting via skip connections enables the transfer of degradation-aware encoder features to the decoder side, thereby enhancing the restoration process. This fundamental correction in the network design gives us significant improvement for all-in-one image restoration tasks, as detailed in the experimental section (Sec. 5).
3.1.2 Architectural details.
At each encoder-decoder level, by simply changing the module weights reuse frequency ($f$), we obtain DyNet’s bulky and lightweight variants. Fig. 2 shows the core architecture of our DyNet. At each encoder-decoder level, we initialize weights ($w^{1}$) for the first transformer block [52] and reuse them for the subsequent blocks. We vary the reuse frequency at each encoder-decoder level and obtain our bulkier and lightweight variants. DyNet comprises four encoder-decoder levels, with the bulky variant (DyNet-L) employing reuse frequencies of $f=[4,6,6,8]$ at encoder-decoder level 1 to level 4, respectively. In contrast, the lightweight variant (DyNet-S) applies weights reuse frequencies $f=[2,3,3,4]$ for encoder-decoder level 1 to level 4, respectively. Further, we incorporate a prompt block [36] at each skip connection, which enhances the encoder features being transferred to the decoder from the encoder. Overall, there are three prompt blocks deployed at each level, each consisting of five prompt components.
Figure 3: The proposed dynamic pre-training strategy is shown. Given a clean image, we create a degraded version by injecting noise (Gaussian or Random), JPEG artifacts and random masking within the same image. Two variants (small and large) of our DyNet are then trained in parallel branches to reconstruct the clean image from masked degraded inputs. Notably, the weights are shared between both variants (DyNet-S and DyNet-L) since they are based on the same architecture but with an intra-network weight-sharing scheme with varying frequencies of block repetition. One of the two parallel branches is randomly activated in a single forward pass of the model (Flag being the binary indicator variable to show branch activation). The dotted lines show the network path is inactivated (Flag=0). An L1 loss is used to calculate pixel differences between reconstructed outputs and targets to update the shared weights of both branches. Both the inter and intra-model weight-sharing as well as random activation of branches lead to a significant reduction in GPU hours required for pre-training.
3.2 Dynamic Pre-training Strategy
In recent times, large-scale pre-training has become a key strategy to improve given network performance. Initializing the network with pre-trained weights provides a robust foundation, boosting performance even with fewer parameters. However, this strategy is resource-intensive, demanding considerable computational power and GPU hours. Therefore, we propose a dynamic pre-training that is capable of simultaneous training of multiple variants of the network. These networks are unified by shared network modules weight but varying network depth. This strategy enables the concurrent training of a diverse set of models tailored to various computational needs and task complexities, all while leveraging a shared baseline architecture. Through the proposed dynamic pre-training strategy, we train both DyNet-L and DyNet-S versions of our DyNet concurrently in a single training session, achieving a 50% reduction in GPU hours.
DyNet-L and DyNet-S utilize the same underlying weights for the transformer blocks initialized at each encoder-decoder level. They just differ in the reuse frequency of these initialized transformer blocks at each encoder-decoder level. Thus, during training iterations, we randomly alternate between DyNet-L and DyNet-S ensuring the optimization of the shared underlying weights as shown in Fig. 3. Moreover, we improve the generalization capability of our DyNet variants by adopting an input masking strategy akin to that proposed in masked auto-encoders [19]. We randomly mask portions of an image and train the DyNet variants to reconstruct these masked regions in a self-supervised fashion. The dataset used for this training is described below.
4 Million-IRD: A Dataset for Image Restoration
Large-scale pre-training for image restoration effectively demands large-scale, high-quality and high-resolution datasets. The current image restoration datasets [27, 2, 35], when combined, offer only a few thousand images. This is considerably inadequate compared to the extensively large-scale datasets [42, 41] utilized in pre-training for other high-level tasks, such as visual recognition, object detection, and segmentation. The relatively small size of the existing training sets for image restoration restricts the performance capabilities of the underlying networks. More importantly, we are motivated by the scaling laws in high-level tasks that demonstrate the large-scale pre-taining can enable even lightweight, efficient model designs to reach the performance mark of much heavier models [1]. To address this gap, we introduce a new million-scale dataset named Million-IRD having $\sim$2M high-quality, high-resolution images, curated specifically for the pre-training of models for image restoration tasks. Examples from our Million-IRD dataset are shown in the Fig. 4. Our data collection and pre-processing pipeline are discussed below.
4.1 Data collection and Pre-processing
[…]
Figure 4: On the left: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs. This includes a variety of textures, scenes from nature, sports activities, images taken during the day and at night, intricate textures, wildlife, shots captured from both close and distant perspectives, forest scenes, pictures of monuments, etc. On the right: Sample low-quality images filtered out during the data pre-processing phase (Sec. 4.1). These images were excluded due to being blurry, watermarked, predominantly featuring flat regions, representing e-commerce product photos, or being noisy or corrupted from artifacts.
4.2 Data post-processing
Overall, our Million-IRD dataset has 2.09 million high-quality, high-resolution images. Each image in our dataset comes with metadata detailing its download link and the image resolution. A detailed breakdown of images and their respective sources is provided in the supplementary material. We extract high-resolution non-overlapping patches (of spatial size $512^{2}$) from each image, followed by the application of a flat region detector which eliminates any patch comprising more than 50% flat area. This post-processing phase allows us to assemble a pool of $\sim$8 million diverse image patches, tailored for the pre-training phase of our DyNet variants.
5 Experiments
We validate the proposed DyNet across three key image restoration tasks: dehazing, deraining, and denoising. In line with the existing PromptIR [36], our experiments are conducted under two distinct setups: (a) All-in-One, wherein a single model is trained to handle all three degradation types, and (b) Single-task, where individual models are trained for each specific image restoration task.
5.1 Implementation Details
Dynamic Pre-training. For robust weights initialization, we conduct a dynamic pre-training for two variants of our DyNet, i.e., DyNet-L and DyNet-S. Both variants share the same underline weights but differ in their transformer block reuse frequency at each encoder-decoder level. For the encoder-decoder levels 1 to 4, we set transformer block weights reuse frequencies of [4,6,6,8] for DyNet-L and [2,3,3,4] for DyNet-S. We use all $\sim$8M patches of size 5122 from our Million-IRD dataset for the dynamic pre-training. We randomly crop a 1282 region from each patch, employing a batch size of 32. Each of these cropped patches undergoes a random augmentation to 50% of its area using either JPEG compression, Gaussian noise, or random noise with varying levels. Furthermore, we mask 30% of the patch, leaving the remaining portion unaltered. In total, 80% of the input patch undergoes modification, either through degradation or masking. For weight optimization, we use the L1 loss and Adam optimizer with parameters $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999, setting the learning rate to 1e-4 for 1 million iterations. Using the setup described above, we simultaneously pre-train both DyNet-L and DyNet-S. At any given iteration, we randomly switch between the two variants. Intriguingly, with each iteration, the shared underlying weights are optimized as shown in Fig. 3. As a result, by the end of this single pre-training session, we get DyNet-L and DyNet-S sharing the same trained underlying weights but differ in network depth, making them suitable for various challenges, including robustness and efficiency.
Datasets. For all-in-one and single-task implementations, we follow the same training protocol as existing PromptIR [36] and finetune the pre-trained DyNet-S and DyNet-L. Following [26, 36], we prepare datasets for various restoration tasks. Specifically, for image denoising, we combine the BSD400 [5] (400 training images) and WED [29] (4,744 images) datasets, adding Gaussian noise at levels $\sigma\in{15,25,50}$ to create noisy images. Testing is conducted on the BSD68 [30] and Urban100 [20] datasets. For deraining, we use Rain100L [50] with 200 training and 100 testing clean-rainy image pairs. Dehazing employs the SOTS [25] dataset, featuring 72,135 training and 500 testing images. To develop a unified model for all tasks, we merge these datasets, fine-tune pre-trained DyNet-L and DyNet-S for 120 epochs and directly evaluate them across different tasks. While, for a single task, the pre-trained DyNet-L is fine-tunned for 120 epochs on a respective training set of the task.
Table 1: Comparison results in the All-in-one restoration setting. Our DyNet-L model outperforms PromptIR [36] by 0.82 dB on average across tasks. Additionally, our DyNet-S model achieves a 0.59 dB average improvement over PromptIR [36], with reductions of 31.34% in parameters and 56.75% in GFlops.
Comparative Dehazing Deraining Denoising on BSD68 dataset [30]) Average
Methods on SOTS [25] on Rain100L [16] $\sigma=15$ $\sigma=25$ $\sigma=50$
BRDNet [44] 23.23/0.895 27.42/0.895 32.26/0.898 29.76/0.836 26.34/0.836 27.80/0.843
LPNet [17] 20.84/0.828 24.88/0.784 26.47/0.778 24.77/0.748 21.26/0.552 23.64/0.738
FDGAN [12] 24.71/0.924 29.89/0.933 30.25/0.910 28.81/0.868 26.43/0.776 28.02/0.883
MPRNet [54] 25.28/0.954 33.57/0.954 33.54/0.927 30.89/0.880 27.56/0.779 30.17/0.899
DL[16] 26.92/0.391 32.62/0.931 33.05/0.914 30.41/0.861 26.90/0.740 29.98/0.875
AirNet [26] 27.94/0.962 34.90/0.967 33.92/0.933 31.26/0.888 28.00/0.797 31.20/0.910
PromptIR [36] 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913
DyNet-S (Ours) 31.51/0.980 38.11/0.981 34.06/0.935 31.41/0.891 28.15/0.802 32.65/0.918
DyNet-L (Ours) 31.98/0.981 38.71/0.983 34.11/0.936 31.44/0.892 28.18/0.803 32.88/0.920
Figure 5: Comparative analysis of image denoising by all-in-one methods on the BSD68 dataset [30]. Our DyNet-L reduces noise, producing more sharp and clear image compared to the PromptIR [36].
5.2 Comparisons on multiple degradations under All-in-one setting
We benchmark our DyNet against a range of general image restoration models [44, 17, 12, 54] and specific all-in-one solutions [16, 26, 36], as shown in Table 1. On average, across various tasks, DyNet-L and DyNet-S outperform the previously best PromptIR [36] by $0.82$ dB and $0.59$ dB, respectively, with DyNet-S also cutting down on 56.75% of parameters and 31.34% of GFlops. For image denoising, DyNet-S achieves a $1.09$ dB higher average PSNR compared to DL[16], with Fig. 5 showcasing its ability to deliver noise-free images with improved structural integrity. Notably, DyNet-L sets a new benchmark in image deraining and dehazing with improvements of $2.34$ dB and $1.4$ dB in PSNR, respectively. Visual comparisons in Fig. 6 and  7 demonstrate DyNet’s superior capability in removing rain and haze, respectively, producing more cleaner images than PromptIR [36].
Figure 6: Comparative analysis of image deraining by all-in-one methods on the Rain100L dataset [16]. Our DyNet-L successfully eliminates rain streaks, producing clear, rain-free images as compared to the recent PromptIR [36].
Figure 7: Comparative analysis of image dehazing by all-in-one methods on the SOTS dataset [25]. Our approach reduces haze, producing more clear image compared to the PromptIR [36].
Table 2: Dehazing results in the single-task setting on the SOTS benchmark dataset [25]. Our DyNet-L achieves a boost of $0.76$ dB over PromptIR [26].
DehazeNet[6] MSCNN[39] AODNet[24] EPDN[37] FDGAN[12] AirNet[26] Restormer[52] PromptIR [36] DyNet-L
22.46/0.851 22.06/0.908 20.29/0.877 22.57/0.863 23.15/0.921 23.18/0.900 30.87/0.969 31.31/0.973 32.07/0.982
5.3 Comparisons on single degradation
We assess DyNet-L’s efficacy in single-task settings, where distinct models are tailored to specific restoration tasks, demonstrating the effectiveness of content-adaptive prompting through prompt blocks. Our results, as shown in Table 2, indicate DyNet-L’s superior performance with an improvement of $0.76$ dB over PromptIR [36] and $1.2$ dB over Restormer [52] in dehazing. This pattern is consistent across other tasks, including deraining and denoising. Notably, compared to PromptIR [36], DyNet-L achieves performance gains of $1.81$ dB in deraining (refer to Table 3) and achieves a $0.13$ dB improvement in denoising at a noise level of $\sigma=50$ on the Urban100 dataset (Table 4).
Table 3: Deraining results in the single-task setting on Rain100L [16]. Compared to the PromptIR [36], our method yields $1.68$ dB PSNR improvement.
DIDMDN[55] UMR[51] SIRR[48] MSPFN[21] LPNet[17] AirNet[26] Restormer[52] PromptIR DyNet-L
23.79/0.773 32.39/0.921 32.37/0.926 33.50/0.948 33.61/0.958 34.90/0.977 36.74/0.978 37.04/0.979 38.85/0.984
Table 4: Denoising comparisons in the single-task setting on BSD68 [30] and Urban100 [20] datasets. For the challenging noise level of $\sigma=50$ on Urban100 [20], our DyNet-L obtains $0.64$ dB gain compared to AirNet [26].
Comparative BSD68 [30] Urban100 [20]
Methods $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$
CBM3D [10] 33.50/0.922 30.69/0.868 27.36/0.763 33.93/0.941 31.36/0.909 27.93/0.840
DnCNN [57] 33.89/0.930 31.23/0.883 27.92/0.789 32.98/0.931 30.81/0.902 27.59/0.833
IRCNN [58] 33.87/0.929 31.18/0.882 27.88/0.790 27.59/0.833 31.20/0.909 27.70/0.840
FFDNet [59] 33.87/0.929 31.21/0.882 27.96/0.789 33.83/0.942 31.40/0.912 28.05/0.848
BRDNet [44] 34.10/0.929 31.43/0.885 28.16/0.794 34.42/0.946 31.99/0.919 28.56/0.858
AirNet [26] 34.14/0.936 31.48/0.893 28.23/0.806 34.40/0.949 32.10/0.924 28.88/0.871
Restormer [52] 34.29/0.937 31.64/0.895 28.41/0.810 34.67/0.969 32.41/0.927 29.31/0.878
PromptIR [36] 34.34/0.938 31.71/0.897 28.49/0.813 34.77/0.952 32.49/0.929 29.39/0.881
DyNet-L (Ours) 34.40/0.938 31.76/0.897 28.53/0.813 34.84/0.952 32.60/0.930 29.52/0.882
Table 5: Ablation experiments for different variants of the proposed DyNet.
Comparative Dehazing Deraining Denoising on BSD68 dataset [30]) Average Par GFlops
Methods on SOTS  [25] on Rain100L [16] $\sigma=15$ $\sigma=25$ $\sigma=50$
PromptIR [36] 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913 37M 242.355
(a) Results without masked dynamic pre-training on Million-IRD
DyNet-S 30.10/0.976 37.10/0.975 33.94/0.931 31.28/0.880 28.03/0.789 32.09/0.910 16M 166.38
DyNet-L 30.67/0.977 37.68/0.975 33.97/0.933 31.31/0.889 28.04/0.797 32.33/0.914 16M 242.35
(b) Results with masked dynamic pre-training on Million-IRD
DyNet-S 31.51/0.980 38.11/0.981 34.06/0.935 31.41/0.891 28.15/0.802 32.65/0.918 16M 166.38
DyNet-L 31.98/0.981 38.71/0.983 34.11/0.936 31.44/0.892 28.18/0.803 32.88/0.920 16M 242.35
5.4 Ablation studies
To study the impact of different components, we conduct various ablation experiments in the all-in-one setting, as summarized in Table 5. Table 5(a) illustrates that our DyNet-L and DyNet-S achieves comparable performance to the baseline PromptIR [36] with reductions of 56.75% in parameters and 31.34% in GFlops, without the proposed dynamic pre-training on Million-IRD dataset. This shows that the proposed fundamental correction of placing prompt block at skip connections is effective compared to on the decoder side as given in PromptIR [36]. Furthermore, pre-training on our Million-IRD dataset brings a significant improvement, resulting in an average PSNR increase of 0.55 dB for DyNet-L compared to the DyNet-L without pre-training as shown in Table 5(b). A similar effect is observed for our DyNet-S variant. Moreover, our proposed dynamic training strategy reduces GPU hours by 50% when training variants of the DyNet. Overall, each of our contributions (i.e., implicit prompting via skip connections, dynamic network architecture, Million-IRD dataset, and dynamic pre-training strategy) synergistically enhances the performance of the proposed approach, while significantly reducing GFlops and learning parameters compared to the baseline PromptIR [36].
Training model with different combinations of degradation. In Table 6, we compare the performance of our DyNet-S in all-in-one setting on aggregated datasets, assessing how varying combinations of degradation types affect it’s effectiveness. All the models in this ablation study are trained for 80 epochs. Here in Table 6, we assess how varying combinations of degradation types (tasks) influence the performance of our DyNet. Notably, the DyNet trained for a combination of the derain and de-noise tasks exhibits better performance for image deraining than the DyNet trained on the combination of derain and dehaze. This shows that some degradations are more relevant for the model to benefit from when trained in a joint manner.
Table 6: Performance of the proposed DyNet-S on different combinations of degradation types i.e., removal of noise, rain, and haze.
Degradation Denoising on BSD68 dataset [30] Deraining on Dehazing on
Noise Rain Haze $\sigma=15$ $\sigma=25$ $\sigma=50$ Rain100L [16] SOTS [25]
34.30/0.938 31.65/0.897 28.42/0.813 - -
- - - 38.72/0.984 -
- - - - 31.45/0.980
34.19/0.936 31.54/0.894 28.30/0.808 38.61/0.984 -
34.11/0.935 31.45/0.893 28.18/0.802 - 31.69/0.981
- - - 38.16/0.982 31.55/0.980
34.06/0.935 31.41/0.891 28.15/0.802 38.11/0.981 31.51/0.980
6 Conclusion
This paper introduces a novel weight-sharing mechanism within the Dynamic Network (DyNet) for efficient all-in-one image restoration tasks, significantly improving computational efficiency with a performance boost. By adjusting module weight-reuse frequency, DyNet allows for seamless alternation between bulkier and lightweight models. The proposed dynamic pre-training strategy simultaneously trains bulkier and lightweight models in a single training session. Thus saving 50% GPU hours compared to traditional training strategy. The ablation study shows that the accuracy of both bulky and lightweight models significantly boosts with the proposed dynamic large-scale pre-training on our Million-IRD dataset. Overall, our DyNet significantly improves the all-in-one image restoration performance with a 31.34% reduction in GFlops and a 56.75% reduction in parameters compared to the baseline models.
Supplemental Material
Here, we have discussed the details about the transformer block used in our DyNet, additional sample images from our Million-IRD dataset, additional visual results comparison between PromptIR [36] and the proposed DyNet-L in all-in-one setting.
Figure S1: Overview of the Transformer block used in our DyNet network. The Transformer block is composed of two sub-modules, the Multi Dconv head transposed attention module (MDTA) and the Gated Dconv feed-forward network (GDFN).
A Transformer Block Details
As described in Section 3.1 of the main manuscript, here, we have discussed the transformer block utilized within our DyNet network architecture, detailing its sub-modules multi deconv head transposed attention (MDTA) and gated deconv feed-forward network (GDFN). Initially, input features, represented as $R\in H_{l}\times W_{l}\times C_{l}$, are processed through the MDTA module. Within this, Layer normalization is used to normalize the input features. This is followed by the application of $1\times 1$ convolutions and then $3\times 3$ depth-wise convolutions, which serve to transform the features into Query (Q), Key (K), and Value (V) tensors. A key feature of the MDTA module is its focus on calculating attention across the channel dimensions, instead of the spatial dimensions, which significantly reduces computational demands. For channel-wise attention, the Q and K tensors are reshaped from $H_{l}\times W_{l}\times C_{l}$ to $H_{l}W_{l}\times C_{l}$ and $C_{l}\times H_{l}W_{l}$ dimensions, respectively. This reshaping facilitates the computation of the dot product and leads to a transposed attention map of $C_{l}\times C_{l}$ dimensions. This process incorporates bias-free convolutions and executes attention computation simultaneously across multiple heads. Following the MDTA Module, the features undergo further processing in the GDFN module. Within this module, the input features are initially expanded by a factor of $\gamma$ through the use of $1\times 1$ convolutions. Subsequently, these expanded features are processed with $3\times 3$ convolutions. This procedure is executed along two parallel pathways. The output from one of these pathways is subjected to activation through the GeLU non-linear function. The activated feature map is then merged with the output from the alternative pathway via an element-wise multiplication.
B Visual Results Comparison
We show additional visual results comparison between PromptIR [36] and our DyNet-L under from all-in-one setting.
Figure S2: Comparative analysis of image denoising by all-in-one methods on the BSD68 dataset [30]. Our DyNet-L reduces noise, producing more sharp and clear image compared to the PromptIR [36].
Figure S3: Comparative analysis of image deraining by all-in-one methods on the Rain100L dataset [16]. Our DyNet-L successfully eliminates rain streaks, producing clear, rain-free images as compared to the recent PromptIR [36].
Figure S4: Comparative analysis of image dehazing by all-in-one methods on the SOTS dataset [25]. Our approach reduces haze, producing more clear image compared to the PromptIR [36].
C Breakdown of images in Our Million-IRD Dataset
We combine the existing high-quality, high-resolution natural image datasets such as LSDIR [27], DIV2K [2], Flickr2K [35], and NTIRE [4]. Collectively these datasets have 90K images having spatial size ranging between $<1024^{2},4096^{2}>$. Their breakdown is given in Table S1. We also show the additional sample images from our Million-IRD dataset in Fig. S5 and Fig. S6.
Table S1: Overview of the number of images curated from each database for our Million-IRD dataset.
Dataset NTIRE [4] DIV2K [2] Flikr2K [35] LSDIR [27] Laion-HR [42]
#Images 56 2,000 2,000 84,991 2M
Figure S5: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs. This includes a variety of textures, scenes from nature, sports activities, images taken during the day and at night, intricate textures, wildlife, shots captured from both close and distant perspectives, forest scenes, pictures of monuments, etc.
Figure S6: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs. This includes a variety of textures, scenes from nature, sports activities, images taken during the day and at night, intricate textures, wildlife, shots captured from both close and distant perspectives, forest scenes, pictures of monuments, etc.

### v2

@@
 Further, to establish robust weights initialization, we introduce a dynamic pre-training strategy that trains variants of the proposed DyNet concurrently, thereby achieving a 50% reduction in GPU hours.
-To tackle the unavailability of large-scale dataset required in pre-training, we curate a high-quality, high-resolution image dataset named Million-IRD having 2M image samples.
+Our dynamic pre-training strategy eliminates the need for maintaining separate checkpoints for each variant, as all models share a common set of checkpoints, varying only in model depth.
+This efficient strategy significantly reduces storage overhead and enhances adaptability.
+To tackle the unavailability of large-scale dataset required in pre-training, we curate a high-quality, high-resolution image dataset named Million-IRD, having 2M image samples.
 We validate our DyNet for image denoising, deraining, and dehazing in all-in-one setting, achieving state-of-the-art results with 31.34% reduction in GFlops and a 56.75% reduction in parameters compared to baseline models.
 The source codes and trained models are available at https://github.com/akshaydudhane16/DyNet.
-Keywords:
-All-in-one Image Restoration Foundation model
 1 Introduction
@@
 Despite several advancements in IR, diverse degradation types and severity levels present in images continue to pose a significant challenge.
-The majority of existing methods [46, 31, 13, 8, 15, 52, 3, 14, 47] learn image priors implicitly, requiring separate network training for different degradation types, levels, and datasets.
+The majority of existing methods (Tu et al., 2022; Zamir et al., 2022; Wang et al., 2021; Zamir et al., 2021; 2020b; Chen et al., 2022) learn image priors implicitly, requiring separate network training for different degradation types, levels, and datasets.
 Further, these methods require a prior knowledge of image degradation for effective model selection during testing, thereby lacking generality to cater diverse degradations.
 All-in-one restoration aims to restore images with an unknown degradation using a single unified model.
-Recent advancements, such as AirNet [26] and PromptIR [36], have addressed the all-in-one restoration challenge by employing contrastive learning and implicit visual prompting techniques, respectively.
-Specifically, the state-of-the-art PromptIR [36] employs an implicit prompting to learn degradation-aware prompts on the decoder side, aiming to refine the decoder features.
+Recent advancements, such as AirNet (Li et al., 2022) and PromptIR (Potlapalli et al., 2023), have addressed the all-in-one restoration challenge by employing contrastive learning and implicit visual prompting techniques, respectively.
+Specifically, the state-of-the-art PromptIR (Potlapalli et al., 2023) employs an implicit prompting to learn degradation-aware prompts on the decoder side, aiming to refine the decoder features.
 This approach, while intriguing, does not extend to the refinement of encoder features, which are left unprocessed.
@@
 The practical application of such models is therefore challenging due to the high computational requirements, especially in resource-constraint hand-held devices.
-However, choosing lightweight models invariably leads to a trade-off between accuracy and efficiency [56, 28].
-Most of these models aim to reduce parameter counts by strided convolutions and feature channel splitting [23, 28, 22].
-Some strategies also operate within the frequency domain to reduce the computational demands linked to attention mechanisms [62] or non-local operations [18] while few methods like [9, 49] partition the feature space for efficient processing, or split the attention across dimensions to improve the computational efficiency [61, 9].
-Towards optimizing all-in-one IR efficiency without sacrificing performance, in this paper, we introduce a novel weights-sharing mechanism.
+However, choosing lightweight models invariably leads to a trade-off between accuracy and efficiency (Zhang et al., 2020a; Li et al., 2023b).
+Most of these models aim to reduce parameter counts by strided convolutions and feature channel splitting (Kong et al., 2022; Li et al., 2023b; Kligvasser et al., 2018).
+Some strategies also operate within the frequency domain to reduce the computational demands linked to attention mechanisms (Zhou et al., 2023) or non-local operations (Guo et al., 2021) while few methods like (Cui & Knoll, 2023; Xie et al., 2021) partition the feature space for efficient processing, or split the attention across dimensions to improve the computational efficiency (Zhao et al., 2023; Cui & Knoll, 2023).
+To optimize all-in-one IR efficiency without sacrificing performance, in this paper, we propose a novel weights-sharing mechanism.
 In this scheme, the weights of a network module are shared with its subsequent modules in a series.
@@
 We deploy the proposed weights-sharing mechanism in an encoder-decoder style architecture named Dynamic Network (DyNet).
-Apart from the computational efficiency, the proposed DyNet provides exceptional flexibility as by merely changing the module weights reuse frequency, one can easily adjust the network depth and correspondingly switch between its bulkier and light-weight variants during training.
+Apart from the computational efficiency, the proposed DyNet provides exceptional flexibility as by merely changing the module weights reuse frequency; one can easily adjust the network depth and correspondingly switch between its bulkier and lightweight variants.
+Model variants share the same checkpoints, differing only in depth, eliminating the need for separate checkpoints and saving disk space.
 While DyNet offers great adaptability with orders of speedup, a common challenge with lightweight networks is the potential compromise on overall accuracy.
@@
 However, large-scale pre-training is computationally intensive and requires significant GPU hours.
-Figure 1: Left: At any given encoder-decoder level, (a) Series of transformer blocks in existing PromptIR [36], (b) The proposed transformer block weights sharing mechanism in our DyNet-L.
-We initialize one transformer block at each encoder-decoder level and shared its weights with the subsequent $N$ transformer blocks.
-(c) Our lightweight variant DyNet-S.
-Here, the number of transformer blocks are halved and weights of the first transformer block are shared with $N/2$ subsequent transformer blocks.
-Right: A plot of average PSNR in All-in-one restoration setting vs GFlops and parameters (in millions).
-The proposed DyNet-S achieves boost of 0.59 dB with 31.34% reduction in GFlops and a 56.75% reduction in parameters compared to the baseline PromptIR [36].
+Figure 1: Left: At any given encoder-decoder level: (a) Transformer blocks in PromptIR, (b) and (c) Our DyNet-L and DyNet-S use the proposed weights sharing mechanism, initializing one transformer block and sharing its weights with subsequent blocks.
+Right: A plot of average PSNR in All-in-one IR setting vs GFlops and parameters (in millions).
+Our DyNet-S boosts performance by 0.43 dB, while reducing GFlops by 31.34% and parameters by 56.75% compared to PromptIR.
 Therefore, we introduce an efficient and potent dynamic pre-training strategy capable of training both bulkier and lightweight network variants within a single pre-training session, thereby achieving significant savings of 50% in GPU hours.
@@
 From this dataset, we extract 8 million non-overlapping high-resolution patches, each sized 512$\times$512, utilized for the pre-training of variants of the proposed DyNet.
-Through the synergy of these techniques, our proposed DyNet achieves an average gain of 0.82 dB for image de-noising, deraining, and dehazing within an all-in-one setting, with 31.34% reduction in GFlops and a 56.75% reduction in network parameters compared to the baseline; refer Fig.
+Through the synergy of these techniques, our proposed DyNet achieves an average gain of 0.68 dB for image de-noising, deraining, and dehazing within an all-in-one setting, with 31.34% reduction in GFlops and a 56.75% reduction in network parameters compared to the baseline; refer Fig.
 1.
 2 Related Work
-Image restoration seeks to restore images from their degraded low-quality versions, with the restoration process varying considerably across different tasks.
-Research in image restoration has predominantly concentrated on addressing single degradation challenges [52, 40, 11, 53, 38, 58, 45, 34, 60].
+Image restoration (IR) seeks to restore images from their degraded low-quality versions, with the restoration process varying considerably across different tasks.
+Research in IR has predominantly concentrated on addressing single degradation challenges (Zamir et al., 2022; Ren et al., 2020; Dong et al., 2020a; Zamir et al., 2020a; Ren et al., 2021; Zhang et al., 2017b; Tsai et al., 2022; Nah et al., 2022; Zhang et al., 2020b).
 Although significant progress has been made in single degradation restoration techniques, research in multi-degradation restoration using a unified model is still relatively under-explored.
 Multi-degradation restoration, however, provides more practical advantages as the information about input degradation type is not required to select the task-specific model and it also enhances model storage efficiency over single-degradation restoration.
-The critical challenge in multi-task image restoration is developing a single model capable of addressing various types of degradation and precisely restoring low-quality images.
-The IPT model [7], for instance, relies on prior knowledge of the corruption affecting the input image, utilizing a pre-trained transformer backbone equipped with distinct encoders and decoder heads tailored for multi-task restoration.
-However, in blind image restoration, we do not have such prior information about the degradation.
-In this direction, Li et al.
-[26] propose a unified network for multiple tasks such as denoising, deraining, and dehazing, employing an image encoder refined through contrastive learning.
+The critical challenge in multi-task IR is developing a single model capable of addressing various types of degradation and precisely restoring low-quality images.
+The IPT model (Chen et al., 2021), for instance, relies on prior knowledge of the corruption affecting the input image, utilizing a pre-trained transformer backbone equipped with distinct encoders and decoder heads tailored for multi-task restoration.
+However, in blind IR, we do not have such prior information about the degradation.
+In this direction,  (Li et al., 2022) propose a unified network for multiple tasks such as denoising, deraining, and dehazing, employing an image encoder refined through contrastive learning.
 However, it requires a two-stage training process, where the outcome of contrastive learning relies on selecting positive and negative pairs and the availability of large data samples.
-Further, motivated by easy implementation and broad applicability of prompts, PromptIR [36] introduces a concept where degradation-aware prompts are learned implicitly within an encoder-decoder architecture for all-in-one image restoration.
-PromptIR applies implicit prompting to the decoder side to refine the decoder features.
-However, it does not extend to the refinement of encoder features, which are left unprocessed.
-Although the above-discussed methods tackle multi-task image restoration, they suffer from low efficiency due to a large number of parameters and substantial GFlops consumption.
+Further, motivated by easy implementation and broad applicability of prompts, PromptIR (Potlapalli et al., 2023) introduces a concept where degradation-aware prompts are learned implicitly within an encoder-decoder architecture for all-in-one IR.
+PromptIR uses implicit prompting to refine decoder features but does not extend this refinement to the encoder features, leaving them unprocessed.
+Although the above-discussed methods tackle multi-task IR, they suffer from low efficiency due to a large number of parameters and substantial GFlops consumption.
 Consequently, deploying these models in environments where efficiency is crucial factor becomes difficult due to their demanding computational needs.
-However, opting for lightweight models introduces an inevitable accuracy-efficiency trade-off [56].
-The primary cause is that many current methods focus on reducing parameter counts through techniques like strided convolutions and feature channel splitting [23, 28, 22].
-Additionally, few approaches work within the frequency domain to reduce the computational burden associated with attention mechanisms [62] or non-local operations [18].
-Furthermore, methods like [9, 49] partition the feature space for efficient processing, while [61, 9] split the attention across dimensions to improve the computational efficiency.
-In contrast, our proposed approach focuses on an efficient and scalable all-in-one image restoration model that incorporates a module weights-sharing mechanism.
+However, opting for lightweight models introduces an inevitable accuracy-efficiency trade-off (Zhang et al., 2020a).
+The primary cause is that existing methods (Kong et al., 2022; Li et al., 2023b; Kligvasser et al., 2018) focus on reducing parameter counts through techniques like strided convolutions and feature channel splitting.
+Further, few approaches work in frequency domain to reduce the computational burden associated with attention mechanisms (Zhou et al., 2023) or non-local operations (Guo et al., 2021).
+Furthermore, methods like (Cui & Knoll, 2023; Xie et al., 2021) partition the feature space for efficient processing, while (Zhao et al., 2023; Cui & Knoll, 2023) split the attention across dimensions to improve the computational efficiency.
+In contrast, our approach focuses on an efficient and scalable all-in-one image restoration model incorporating a weights-sharing mechanism.
 We demonstrate that strategic fundamental adjustments to the network architecture result in substantial performance improvements.
-Figure 2: The overall pipeline of the proposed Dynamic Network (DyNet) for all-in-one image restoration.
-DyNet enhances a low-quality input image using a four-level encoder-decoder architecture.
-A distinctive aspect of DyNet lies in its weight-sharing strategy: at each level, the initial transformer block’s weights are sequentially passed to subsequent blocks, significantly reducing the network’s parameters and enhancing its flexibility.
-This approach allows for easy modification of DyNet’s complexity, toggling between more bulkier or lightweight variants by adjusting the frequency of weight sharing among the blocks at each encoder-decoder level.
-Moreover, we maintain the encoder-decoder feature consistency by implicitly learning degradation-aware prompts at skip connections rather than on the decoder side as in PromptIR [36].
+Figure 2: The proposed Dynamic Network (DyNet) pipeline for all-in-one image restoration.
+DyNet enhances a low-quality input image using a 4-level encoder-decoder architecture.
+A distinctive aspect of DyNet lies in its weight-sharing strategy.
+At each level, the initial transformer block’s weights are shared with the subsequent blocks, significantly reducing the network parameters and enhancing its flexibility.
+This approach allows for easy adjustment of DyNet’s complexity, switching between large and small variants by modifying the frequency of weight sharing across the encoder-decoder blocks.
+Moreover, we maintain the encoder-decoder feature consistency by implicitly learning degradation-aware prompts at skip connections rather than on the decoder side as in PromptIR.
 3 Proposed Approach
-Existing state-of-the-art all-in-one image restoration (IR) approaches incur a high computational footprint and do not provide the flexibility to support models with variable depths during training.
+Existing SoTA all-in-one image restoration (IR) approaches offer a high computational footprint and do not provide the flexibility to change model complexity ”on-the-go” during deployment.
 Towards more accurate, but lightweight and flexible architecture design for all-in-one IR, this work proposes a Dynamic Network Architecture (Sec.
@@
 3.1 Dynamic Network (DyNet) Architecture
-Our lightweight architecture is based on a novel weight-sharing mechanism for IR models.
+Our network architecture is based on a novel weight-sharing mechanism for IR models.
 This mechanism allows the network module’s weights to be reused across subsequent modules in sequence, thereby significantly reducing the total number of parameters and leading to a more efficient network structure.
@@
 3.1.1 Overall Pipeline.
-The overall pipeline of the proposed DyNet is shown in Fig.
+The proposed DyNet pipeline is shown in Fig.
 2.
@@
 Subsequently, the feature embedding undergoes a 4-level hierarchical encoder-decoder, gradually transforming into deep latent features $\mathbf{F}_{l}$in $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$.
-We utilize the existing Restormer’s [52] Transformer block as a fundamental feature extraction module.
+We utilize the existing Restormer’s (Zamir et al., 2022) Transformer block as a fundamental feature extraction module.
 For every level of the encoder and decoder, we initialize the weights for a first transformer block, which are subsequently reused across the subsequent blocks in a series up to the specified frequency at that level.
@@
 […]
-We gradually increase the reuse frequency of the transformer block from the top level to the bottom level, thereby increasing the network depth.
-The multi-level encoder systematically reduces spatial resolution while increasing channel capacity, facilitating the progressive extraction of low-resolution latent features from a high-resolution input image.
+We gradually increase the reuse frequency of the transformer block from top level to the bottom level, thereby increasing the network depth.
+The multi-level encoder systematically reduces spatial resolution while increasing channel capacity, enabling the extraction of low-resolution latent features from a high-resolution input image.
 Subsequently, the multi-level decoder gradually restores the high-resolution output from these low-resolution latent features.
-To enhance the decoding process, we integrate prompt blocks [36], which learn degradation-aware prompts implicitly through the prompt generation module (PGM) followed by the prompt-interaction module (PIM).
-Unlike the existing PromptIR [36], our approach incorporates degradation-aware implicit prompt blocks at skip connections.
-This implicit prompting via skip connections enables the transfer of degradation-aware encoder features to the decoder side, thereby enhancing the restoration process.
+To enhance the decoding process, we integrate prompt blocks from PromptIR (Potlapalli et al., 2023), which learn degradation-aware prompts implicitly through the prompt generation module (PGM) followed by the prompt-interaction module (PIM).
+Unlike the existing PromptIR, our approach incorporates degradation-aware implicit prompt blocks at skip connections.
+This implicit prompting via skip connections enables the transfer of refined encoder features to the decoder side, thereby enhancing the restoration process.
 This fundamental correction in the network design gives us significant improvement for all-in-one image restoration tasks, as detailed in the experimental section (Sec.
@@
 3.1.2 Architectural details.
-At each encoder-decoder level, by simply changing the module weights reuse frequency ($f$), we obtain DyNet’s bulky and lightweight variants.
+At each encoder-decoder level, by adjusting the module weights reuse frequency ($f$), we obtain DyNet’s bulky and lightweight variants.
 Fig.
 2 shows the core architecture of our DyNet.
-At each encoder-decoder level, we initialize weights ($w^{1}$) for the first transformer block [52] and reuse them for the subsequent blocks.
+At each encoder-decoder level, we initialize weights ($w^{1}$) for the first transformer block and reuse them for the subsequent blocks.
 We vary the reuse frequency at each encoder-decoder level and obtain our bulkier and lightweight variants.
-DyNet comprises four encoder-decoder levels, with the bulky variant (DyNet-L) employing reuse frequencies of $f=[4,6,6,8]$ at encoder-decoder level 1 to level 4, respectively.
-In contrast, the lightweight variant (DyNet-S) applies weights reuse frequencies $f=[2,3,3,4]$ for encoder-decoder level 1 to level 4, respectively.
-Further, we incorporate a prompt block [36] at each skip connection, which enhances the encoder features being transferred to the decoder from the encoder.
-Overall, there are three prompt blocks deployed at each level, each consisting of five prompt components.
+DyNet comprises four encoder-decoder levels, with the larger variant (DyNet-L) employing reuse frequencies of $f=[4,6,6,8]$ at encoder-decoder: level 1 to level 4, respectively.
+In contrast, the smaller variant (DyNet-S) applies weights reuse frequencies $f=[2,3,3,4]$ for encoder-decoder level 1 to level 4, respectively.
+Further, we incorporate a prompt block at each skip connection, which enhances the encoder features being transferred to the decoder from the encoder.
+Overall, there are three prompt blocks: deployed one at each skip connection, each consisting of five prompt components each.
 Figure 3: The proposed dynamic pre-training strategy is shown.
 Given a clean image, we create a degraded version by injecting noise (Gaussian or Random), JPEG artifacts and random masking within the same image.
-Two variants (small and large) of our DyNet are then trained in parallel branches to reconstruct the clean image from masked degraded inputs.
-Notably, the weights are shared between both variants (DyNet-S and DyNet-L) since they are based on the same architecture but with an intra-network weight-sharing scheme with varying frequencies of block repetition.
+Two variants (small and large) of our DyNet are then trained concurrently to reconstruct the clean image from masked degraded inputs.
+Notably, the weights are shared between both variants since they are based on the same architecture but with an intra-network weight-sharing scheme with varying frequencies of block repetition.
 One of the two parallel branches is randomly activated in a single forward pass of the model (Flag being the binary indicator variable to show branch activation).
@@
 An L1 loss is used to calculate pixel differences between reconstructed outputs and targets to update the shared weights of both branches.
-Both the inter and intra-model weight-sharing as well as random activation of branches lead to a significant reduction in GPU hours required for pre-training.
+Both the inter and intra-model weight-sharing and random activation of branches lead to a significant reduction in GPU hours required for pre-training.
 3.2 Dynamic Pre-training Strategy
@@
 However, this strategy is resource-intensive, demanding considerable computational power and GPU hours.
-Therefore, we propose a dynamic pre-training that is capable of simultaneous training of multiple variants of the network.
-These networks are unified by shared network modules weight but varying network depth.
-This strategy enables the concurrent training of a diverse set of models tailored to various computational needs and task complexities, all while leveraging a shared baseline architecture.
-Through the proposed dynamic pre-training strategy, we train both DyNet-L and DyNet-S versions of our DyNet concurrently in a single training session, achieving a 50% reduction in GPU hours.
-DyNet-L and DyNet-S utilize the same underlying weights for the transformer blocks initialized at each encoder-decoder level.
-They just differ in the reuse frequency of these initialized transformer blocks at each encoder-decoder level.
+Therefore, we propose a dynamic pre-training strategy that enables simultaneous training of multiple network variants with shared module weights but varying depths.
+This allows for concurrent training of diverse models tailored to different computational needs and task complexities, all leveraging a shared architecture.
+Through the proposed dynamic pre-training strategy, we train both DyNet-L and DyNet-S variants of our DyNet concurrently in a single training session, achieving a 50% reduction in GPU hours.
+DyNet-L and DyNet-S utilize the same underlying weights for transformer blocks, differing only in the reuse frequency of these blocks at each encoder-decoder level.
 Thus, during training iterations, we randomly alternate between DyNet-L and DyNet-S ensuring the optimization of the shared underlying weights as shown in Fig.
 3.
-Moreover, we improve the generalization capability of our DyNet variants by adopting an input masking strategy akin to that proposed in masked auto-encoders [19].
-We randomly mask portions of an image and train the DyNet variants to reconstruct these masked regions in a self-supervised fashion.
-The dataset used for this training is described below.
+Moreover, we improve the generalization capability of our DyNet variants by adopting an input masking strategy akin to that proposed in masked auto-encoders (He et al., 2022).
+We randomly mask portions of an image and train the DyNet variants to reconstruct masked regions in a self-supervised fashion.
+The training dataset is described below.
 4 Million-IRD: A Dataset for Image Restoration
 Large-scale pre-training for image restoration effectively demands large-scale, high-quality and high-resolution datasets.
-The current image restoration datasets [27, 2, 35], when combined, offer only a few thousand images.
-This is considerably inadequate compared to the extensively large-scale datasets [42, 41] utilized in pre-training for other high-level tasks, such as visual recognition, object detection, and segmentation.
+The current image restoration datasets: LSDIR (Li et al., 2023a), NTIRE (Agustsson & Timofte, 2017), Flickr2K (Online, ), when combined, offer only a few thousand images.
+This is considerably inadequate compared to the extensively large-scale datasets LAION-5B (Schuhmann et al., 2022), ImageNet-21K (Ridnik et al., 2021) utilized in pre-training for other high-level tasks, such as visual recognition, object detection, and segmentation.
 The relatively small size of the existing training sets for image restoration restricts the performance capabilities of the underlying networks.
-More importantly, we are motivated by the scaling laws in high-level tasks that demonstrate the large-scale pre-taining can enable even lightweight, efficient model designs to reach the performance mark of much heavier models [1].
-To address this gap, we introduce a new million-scale dataset named Million-IRD having $\sim$2M high-quality, high-resolution images, curated specifically for the pre-training of models for image restoration tasks.
-Examples from our Million-IRD dataset are shown in the Fig.
-4.
+More importantly, we are motivated by the scaling laws in high-level tasks that demonstrate the large-scale pre-taining can enable even lightweight, efficient model designs to reach the performance mark of much heavier models (Abnar et al., 2021).
+To address this gap, we introduce a new million-scale dataset named Million-IRD having $\sim$2M high-quality, high-resolution images, curated specifically for the pre-training of models for IR tasks.
+Fig.
+4 shows few samples from our Million-IRD dataset.
 Our data collection and pre-processing pipeline are discussed below.
@@
 Each image in our dataset comes with metadata detailing its download link and the image resolution.
-A detailed breakdown of images and their respective sources is provided in the supplementary material.
+A detailed breakdown of images and their respective sources are provided in the Appendix A.
 We extract high-resolution non-overlapping patches (of spatial size $512^{2}$) from each image, followed by the application of a flat region detector which eliminates any patch comprising more than 50% flat area.
@@
 We validate the proposed DyNet across three key image restoration tasks: dehazing, deraining, and denoising.
-In line with the existing PromptIR [36], our experiments are conducted under two distinct setups: (a) All-in-One, wherein a single model is trained to handle all three degradation types, and (b) Single-task, where individual models are trained for each specific image restoration task.
+In line with the existing PromptIR (Potlapalli et al., 2023), our experiments are conducted under two distinct setups: (a) All-in-One, wherein a single model is trained to handle all three degradation types, and (b) Single-task, where individual models are trained for each specific image restoration task.
 5.1 Implementation Details
@@
 Datasets.
-For all-in-one and single-task implementations, we follow the same training protocol as existing PromptIR [36] and finetune the pre-trained DyNet-S and DyNet-L.
-Following [26, 36], we prepare datasets for various restoration tasks.
-Specifically, for image denoising, we combine the BSD400 [5] (400 training images) and WED [29] (4,744 images) datasets, adding Gaussian noise at levels $\sigma\in{15,25,50}$ to create noisy images.
-Testing is conducted on the BSD68 [30] and Urban100 [20] datasets.
-For deraining, we use Rain100L [50] with 200 training and 100 testing clean-rainy image pairs.
-Dehazing employs the SOTS [25] dataset, featuring 72,135 training and 500 testing images.
-To develop a unified model for all tasks, we merge these datasets, fine-tune pre-trained DyNet-L and DyNet-S for 120 epochs and directly evaluate them across different tasks.
-While, for a single task, the pre-trained DyNet-L is fine-tunned for 120 epochs on a respective training set of the task.
+For the all-in-one task implementations, we adopt the training protocol from PromptIR and concurrently fine-tune the pre-trained DyNet-S and DyNet-L in a single session.
+Following PromptIR, we prepare datasets for various restoration tasks.
+Specifically, for image denoising, we combine the BSD400 (Arbelaez et al., 2011) (400 training images) and WED (Ma et al., 2016) (4,744 images) datasets, adding Gaussian noise at levels $\sigma\in[15,25,50]$ to create noisy images.
+Testing is conducted on the BSD68 (Martin et al., 2001) and Urban100 (Huang et al., 2015) datasets.
+For deraining, we use Rain100L (Yang et al., 2020) with 200 training and 100 testing clean-rainy image pairs.
+Dehazing employs the SOTS (Li et al., 2018) dataset, featuring 72,135 training and 500 testing images.
+To build a unified model for all tasks, we merge the datasets and deploy a dynamic finetunning of the pre-trained DyNet for 120 epochs, then evaluate it across multiple tasks.
+The resulting flexible DyNet allows ”on-the-go” adjustments to model depth, switching between DyNet-L and DyNet-S without needing separate model weights.
+Both versions share the same underlying weights, differing only in depth.
+For individual tasks, the pre-trained DyNet-L is fine-tuned for 120 epochs on the corresponding task-specific training set.
 Table 1: Comparison results in the All-in-one restoration setting.
-Our DyNet-L model outperforms PromptIR [36] by 0.82 dB on average across tasks.
-Additionally, our DyNet-S model achieves a 0.59 dB average improvement over PromptIR [36], with reductions of 31.34% in parameters and 56.75% in GFlops.
-Comparative Dehazing Deraining Denoising on BSD68 dataset [30]) Average
-Methods on SOTS [25] on Rain100L [16] $\sigma=15$ $\sigma=25$ $\sigma=50$
-BRDNet [44] 23.23/0.895 27.42/0.895 32.26/0.898 29.76/0.836 26.34/0.836 27.80/0.843
-LPNet [17] 20.84/0.828 24.88/0.784 26.47/0.778 24.77/0.748 21.26/0.552 23.64/0.738
-FDGAN [12] 24.71/0.924 29.89/0.933 30.25/0.910 28.81/0.868 26.43/0.776 28.02/0.883
-MPRNet [54] 25.28/0.954 33.57/0.954 33.54/0.927 30.89/0.880 27.56/0.779 30.17/0.899
-DL[16] 26.92/0.391 32.62/0.931 33.05/0.914 30.41/0.861 26.90/0.740 29.98/0.875
-AirNet [26] 27.94/0.962 34.90/0.967 33.92/0.933 31.26/0.888 28.00/0.797 31.20/0.910
-PromptIR [36] 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913
-DyNet-S (Ours) 31.51/0.980 38.11/0.981 34.06/0.935 31.41/0.891 28.15/0.802 32.65/0.918
-DyNet-L (Ours) 31.98/0.981 38.71/0.983 34.11/0.936 31.44/0.892 28.18/0.803 32.88/0.920
-Figure 5: Comparative analysis of image denoising by all-in-one methods on the BSD68 dataset [30]. Our DyNet-L reduces noise, producing more sharp and clear image compared to the PromptIR [36].
+Our DyNet-L model outperforms PromptIR by 0.68 dB on average across tasks.
+Further, our DyNet-S model achieves a 0.43 dB average improvement over PromptIR, with reductions of 31.34% in parameters and 56.75% in GFlops.
+Comparative Dehazing Deraining Denoising on BSD68 dataset Average
+Methods on SOTS on Rain100L $\sigma=15$ $\sigma=25$ $\sigma=50$ PSNR/SSIM
+BRDNet (Tian et al., 2020) 23.23/0.895 27.42/0.895 32.26/0.898 29.76/0.836 26.34/0.836 27.80/0.843
+LPNet (Gao et al., 2019) 20.84/0.828 24.88/0.784 26.47/0.778 24.77/0.748 21.26/0.552 23.64/0.738
+FDGAN (Dong et al., 2020b) 24.71/0.924 29.89/0.933 30.25/0.910 28.81/0.868 26.43/0.776 28.02/0.883
+MPRNet (Zamir et al., 2021) 25.28/0.954 33.57/0.954 33.54/0.927 30.89/0.880 27.56/0.779 30.17/0.899
+DL (Fan et al., 2019) 26.92/0.391 32.62/0.931 33.05/0.914 30.41/0.861 26.90/0.740 29.98/0.875
+AirNet (Li et al., 2022) 27.94/0.962 34.90/0.967 33.92/0.933 31.26/0.888 28.00/0.797 31.20/0.910
+PromptIR (Potlapalli et al., 2023) 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913
+DyNet-S (Ours) 30.80/0.980 38.02/0.982 34.07/0.935 31.41/0.892 28.15/0.802 32.49/0.918
+DyNet-L (Ours) 31.34/0.980 38.69/0.983 34.09/0.935 31.43/0.892 28.18/0.803 32.74/0.920
+Figure 5: Comparative analysis of image denoising by all-in-one methods on the BSD68 and Urban100 dataset. DyNet reduces noise, produces a sharp and clear images compared to the PromptIR
+Figure 6: Comparative analysis within all-in-one methods for image de-raining on the Rain100L dataset. Our DyNet-L outperforms PromptIR by producing clearer, rain-free images.
+Figure 7: Comparative analysis of image dehazing by all-in-one methods on the real-world hazy images. Our approach reduces haze, producing more clear image compared to the PromptIR.
 5.2 Comparisons on multiple degradations under All-in-one setting
-We benchmark our DyNet against a range of general image restoration models [44, 17, 12, 54] and specific all-in-one solutions [16, 26, 36], as shown in Table 1.
-On average, across various tasks, DyNet-L and DyNet-S outperform the previously best PromptIR [36] by $0.82$ dB and $0.59$ dB, respectively, with DyNet-S also cutting down on 56.75% of parameters and 31.34% of GFlops.
-For image denoising, DyNet-S achieves a $1.09$ dB higher average PSNR compared to DL[16], with Fig.
+We benchmark our DyNet against a range of general IR methods and specific all-in-one solutions, as shown in Table 1.
+On average, across various tasks, DyNet-L and DyNet-S outperform the previously best PromptIR by $0.68$ dB and $0.43$ dB, respectively, and DyNet-S also cuts down 56.75% of parameters and 31.34% of GFlops.
+For image denoising, DyNet-L achieves a $1.11$ dB higher average PSNR compared to DL, and Fig.
 5 showcasing its ability to deliver noise-free images with improved structural integrity.
-Notably, DyNet-L sets a new benchmark in image deraining and dehazing with improvements of $2.34$ dB and $1.4$ dB in PSNR, respectively.
-Visual comparisons in Fig.
-6 and  7 demonstrate DyNet’s superior capability in removing rain and haze, respectively, producing more cleaner images than PromptIR [36].
-Figure 6: Comparative analysis of image deraining by all-in-one methods on the Rain100L dataset [16].
-Our DyNet-L successfully eliminates rain streaks, producing clear, rain-free images as compared to the recent PromptIR [36].
-Figure 7: Comparative analysis of image dehazing by all-in-one methods on the SOTS dataset [25]. Our approach reduces haze, producing more clear image compared to the PromptIR [36].
-Table 2: Dehazing results in the single-task setting on the SOTS benchmark dataset [25]. Our DyNet-L achieves a boost of $0.76$ dB over PromptIR [26].
-DehazeNet[6] MSCNN[39] AODNet[24] EPDN[37] FDGAN[12] AirNet[26] Restormer[52] PromptIR [36] DyNet-L
-22.46/0.851 22.06/0.908 20.29/0.877 22.57/0.863 23.15/0.921 23.18/0.900 30.87/0.969 31.31/0.973 32.07/0.982
+Notably, DyNet-L sets a new benchmark in image deraining and dehazing with improvements of $2.32$ dB and $0.76$ dB in PSNR, respectively.
+In addition, the visual comparisons in Fig.
+6 demonstrate DyNet’s superior capability to remove rain, producing cleaner images than PromptIR.
+For image dehazing, existing methods such as PromptIR are limited to removing neutral color (gray) haze.
+However, in real-world conditions, haze can appear in different colors, such as light blue during the morning or yellow in smog-filled environments, etc.
+As illustrated in Fig.
+7, PromptIR struggles to restore images with nongray haze due to its rigidity to haze-color types.
+Considering different types of haze (increasing the prompt length) during training could potentially address this issue.
+However, considering all types of haze color is not a practical solution.
+Therefore, to enhance our DyNet’s ability to restore different haze-color images, we first restore the color balance using the existing Gray-World Assumption (GWA) algorithm followed by the dehazing through our network, which effectively recovers the haze-free image.
+We present a visual comparison between PromptIR and our DyNet on a variety of real-world hazy images in Fig.
+7.
+For a fair assessment, we also include results from GWA+PromptIR (PromptIR+).
+The results clearly demonstrate that our DyNet outperforms PromptIR in reducing haze effects.
+The simple plug-in GWA module enables our approach to dehaze real-world hazy images without increasing model complexity.
 5.3 Comparisons on single degradation
 We assess DyNet-L’s efficacy in single-task settings, where distinct models are tailored to specific restoration tasks, demonstrating the effectiveness of content-adaptive prompting through prompt blocks.
-Our results, as shown in Table 2, indicate DyNet-L’s superior performance with an improvement of $0.76$ dB over PromptIR [36] and $1.2$ dB over Restormer [52] in dehazing.
+Our results, as shown in Table 2, indicate DyNet-L’s superior performance with an improvement of $0.76$ dB over PromptIR and $1.2$ dB over Restormer in dehazing.
 This pattern is consistent across other tasks, including deraining and denoising.
-Notably, compared to PromptIR [36], DyNet-L achieves performance gains of $1.81$ dB in deraining (refer to Table 3) and achieves a $0.13$ dB improvement in denoising at a noise level of $\sigma=50$ on the Urban100 dataset (Table 4).
-Table 3: Deraining results in the single-task setting on Rain100L [16]. Compared to the PromptIR [36], our method yields $1.68$ dB PSNR improvement.
-DIDMDN[55] UMR[51] SIRR[48] MSPFN[21] LPNet[17] AirNet[26] Restormer[52] PromptIR DyNet-L
+Notably, compared to PromptIR, DyNet-L achieves performance gains of $1.81$ dB in deraining (refer to Table 3) and achieves a $0.24$ dB improvement in denoising at a noise level of $\sigma=50$ on the Urban100 dataset (Table 4).
+Table 2: Dehazing results in the single-task setting on the SOTS (Li et al., 2018) dataset. Our DyNet-L achieves a boost of $0.76$ dB over PromptIR.
+MSCNN AODNet EPDN FDGAN AirNet Restormer PromptIR DyNet-L
+(Ren et al., 2020) (Cai et al., 2016) (Qu et al., 2019) (Dong et al., 2020b) (Li et al., 2022) (Zamir et al., 2022) (Potlapalli et al., 2023) (Ours)
+22.06/0.908 20.29/0.877 22.57/0.863 23.15/0.921 23.18/0.900 30.87/0.969 31.31/0.973 32.07/0.982
+Table 3: Deraining results in the single-task setting on Rain100L (Yang et al., 2020). Compared to the PromptIR, our method yields 1.81 dB PSNR improvement.
+DIDMDN UMR SIRR MSPFN LPNet AirNet Restormer PromptIR DyNet-L
+(Zhang & Patel, 2018) (Yasarla & Patel, 2019) (Wei et al., 2019) (Jiang et al., 2020) (Gao et al., 2019) (Li et al., 2022) (Zamir et al., 2022) (Potlapalli et al., 2023) (Ours)
 23.79/0.773 32.39/0.921 32.37/0.926 33.50/0.948 33.61/0.958 34.90/0.977 36.74/0.978 37.04/0.979 38.85/0.984
-Table 4: Denoising comparisons in the single-task setting on BSD68 [30] and Urban100 [20] datasets.
-For the challenging noise level of $\sigma=50$ on Urban100 [20], our DyNet-L obtains $0.64$ dB gain compared to AirNet [26].
-Comparative BSD68 [30] Urban100 [20]
-Methods $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$
-CBM3D [10] 33.50/0.922 30.69/0.868 27.36/0.763 33.93/0.941 31.36/0.909 27.93/0.840
-DnCNN [57] 33.89/0.930 31.23/0.883 27.92/0.789 32.98/0.931 30.81/0.902 27.59/0.833
-IRCNN [58] 33.87/0.929 31.18/0.882 27.88/0.790 27.59/0.833 31.20/0.909 27.70/0.840
-FFDNet [59] 33.87/0.929 31.21/0.882 27.96/0.789 33.83/0.942 31.40/0.912 28.05/0.848
-BRDNet [44] 34.10/0.929 31.43/0.885 28.16/0.794 34.42/0.946 31.99/0.919 28.56/0.858
-AirNet [26] 34.14/0.936 31.48/0.893 28.23/0.806 34.40/0.949 32.10/0.924 28.88/0.871
-Restormer [52] 34.29/0.937 31.64/0.895 28.41/0.810 34.67/0.969 32.41/0.927 29.31/0.878
-PromptIR [36] 34.34/0.938 31.71/0.897 28.49/0.813 34.77/0.952 32.49/0.929 29.39/0.881
-DyNet-L (Ours) 34.40/0.938 31.76/0.897 28.53/0.813 34.84/0.952 32.60/0.930 29.52/0.882
+Table 4: Denoising comparisons in the single-task setting on BSD68 and Urban100 datasets. For difficult noise level of $\sigma=50$ on Urban100, our DyNet-L obtains $0.75$ dB gain compared to AirNet.
+Comparative Urban100( (Huang et al., 2015)) BSD68( (Martin et al., 2001)) Average
+Methods $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ PSNR/SSIM
+CBM3D (Dabov et al., 2007) 33.93/0.941 31.36/0.909 27.93/0.840 33.50/0.922 30.69/0.868 27.36/0.763 30.79/0.874
+DnCNN (Zhang et al., 2017a) 32.98/0.931 30.81/0.902 27.59/0.833 33.89/0.930 31.23/0.883 27.92/0.789 30.74/0.878
+IRCNN (Zhang et al., 2017b) 27.59/0.833 31.20/0.909 27.70/0.840 33.87/0.929 31.18/0.882 27.88/0.790 29.90/0.864
+FFDNet (Zhang et al., 2018) 33.83/0.942 31.40/0.912 28.05/0.848 33.87/0.929 31.21/0.882 27.96/0.789 31.05/0.884
+BRDNet (Tian et al., 2020) 34.42/0.946 31.99/0.919 28.56/0.858 34.10/0.929 31.43/0.885 28.16/0.794 31.44/0.889
+AirNet (Li et al., 2022) 34.40/0.949 32.10/0.924 28.88/0.871 34.14/0.936 31.48/0.893 28.23/0.806 31.54/0.897
+Restormer (Zamir et al., 2022) 34.67/0.969 32.41/0.927 29.31/0.878 34.29/0.937 31.64/0.895 28.41/0.810 31.79/0.903
+PromptIR (Potlapalli et al., 2023) 34.77/0.952 32.49/0.929 29.39/0.881 34.34/0.938 31.71/0.897 28.49/0.813 31.87/0.902
+DyNet-L (Ours) 34.91/0.953 32.68/0.930 29.63/0.884 34.34/0.938 31.71/0.896 28.47/0.812 31.96/0.902
 Table 5: Ablation experiments for different variants of the proposed DyNet.
-Comparative Dehazing Deraining Denoising on BSD68 dataset [30]) Average Par GFlops
-Methods on SOTS  [25] on Rain100L [16] $\sigma=15$ $\sigma=25$ $\sigma=50$
-PromptIR [36] 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913 37M 242.355
+Comparative Dehazing Deraining Denoising on BSD68 dataset (Martin et al., 2001) Average Par GFlops
+Methods on SOTS on Rain100L $\sigma=15$ $\sigma=25$ $\sigma=50$
+PromptIR 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913 37M 242.355
 (a) Results without masked dynamic pre-training on Million-IRD
@@
 (b) Results with masked dynamic pre-training on Million-IRD
-DyNet-S 31.51/0.980 38.11/0.981 34.06/0.935 31.41/0.891 28.15/0.802 32.65/0.918 16M 166.38
-DyNet-L 31.98/0.981 38.71/0.983 34.11/0.936 31.44/0.892 28.18/0.803 32.88/0.920 16M 242.35
+DyNet-S 30.80/0.980 38.02/0.982 34.07/0.935 31.41/0.892 28.15/0.802 32.49/0.918 16M 166.38
+DyNet-L 31.34/0.981 38.69/0.983 34.09/0.935 31.43/0.892 28.18/0.803 32.74/0.920 16M 242.35
+Table 6: Performance of the proposed DyNet-S on different combinations of degradation types.
+Degradation Denoising on BSD68 dataset Deraining on Dehazing on
+Noise Rain Haze $\sigma=15$ $\sigma=25$ $\sigma=50$ Rain100L SOTS
+34.28/0.938 31.64/0.897 28.42/0.813 - -
+- - - 38.69/0.983 -
+- - - - 30.72/0.980
+34.18/0.936 31.52/0.893 28.31/0.808 38.54/0.983 -
+34.11/0.935 31.44/0.893 28.18/0.802 - 31.08/0.981
+- - - 38.12/0.982 30.92/0.980
+34.07/0.935 31.41/0.892 28.15/0.802 38.02/0.982 30.80/0.980
 5.4 Ablation studies
 To study the impact of different components, we conduct various ablation experiments in the all-in-one setting, as summarized in Table 5.
-Table 5(a) illustrates that our DyNet-L and DyNet-S achieves comparable performance to the baseline PromptIR [36] with reductions of 56.75% in parameters and 31.34% in GFlops, without the proposed dynamic pre-training on Million-IRD dataset.
-This shows that the proposed fundamental correction of placing prompt block at skip connections is effective compared to on the decoder side as given in PromptIR [36].
-Furthermore, pre-training on our Million-IRD dataset brings a significant improvement, resulting in an average PSNR increase of 0.55 dB for DyNet-L compared to the DyNet-L without pre-training as shown in Table 5(b).
+Table 5(a) illustrates that our DyNet-L and DyNet-S achieves comparable performance to the baseline PromptIR with reductions of 56.75% in parameters and 31.34% in GFlops, without the proposed dynamic pre-training on Million-IRD dataset.
+This shows that the proposed fundamental correction of placing prompt block at skip connections is effective compared to on the decoder side as given in PromptIR.
+Furthermore, pre-training on our Million-IRD dataset brings a significant improvement, resulting in an average PSNR increase of 0.41 dB for DyNet-L compared to the DyNet-L without pre-training as shown in Table 5(b).
 A similar effect is observed for our DyNet-S variant.
 Moreover, our proposed dynamic training strategy reduces GPU hours by 50% when training variants of the DyNet.
-Overall, each of our contributions (i.e., implicit prompting via skip connections, dynamic network architecture, Million-IRD dataset, and dynamic pre-training strategy) synergistically enhances the performance of the proposed approach, while significantly reducing GFlops and learning parameters compared to the baseline PromptIR [36].
+Overall, each of our contributions (i.e., implicit prompting via skip connections, dynamic network architecture, Million-IRD dataset, and dynamic pre-training strategy) synergistically enhances the performance of the proposed approach, while significantly reducing GFlops and learning parameters compared to the baseline PromptIR.
 Training model with different combinations of degradation.
@@
 This shows that some degradations are more relevant for the model to benefit from when trained in a joint manner.
-Table 6: Performance of the proposed DyNet-S on different combinations of degradation types i.e., removal of noise, rain, and haze.
-Degradation Denoising on BSD68 dataset [30] Deraining on Dehazing on
-Noise Rain Haze $\sigma=15$ $\sigma=25$ $\sigma=50$ Rain100L [16] SOTS [25]
-34.30/0.938 31.65/0.897 28.42/0.813 - -
-- - - 38.72/0.984 -
-- - - - 31.45/0.980
-34.19/0.936 31.54/0.894 28.30/0.808 38.61/0.984 -
-34.11/0.935 31.45/0.893 28.18/0.802 - 31.69/0.981
-- - - 38.16/0.982 31.55/0.980
-34.06/0.935 31.41/0.891 28.15/0.802 38.11/0.981 31.51/0.980
 6 Conclusion
@@
 Overall, our DyNet significantly improves the all-in-one image restoration performance with a 31.34% reduction in GFlops and a 56.75% reduction in parameters compared to the baseline models.
-Supplemental Material
-Here, we have discussed the details about the transformer block used in our DyNet, additional sample images from our Million-IRD dataset, additional visual results comparison between PromptIR [36] and the proposed DyNet-L in all-in-one setting.
-Figure S1: Overview of the Transformer block used in our DyNet network.
+Appendix A Appendix (Supplementary Material)
+Here, we have discussed the details about the transformer block used in our DyNet, additional sample images from our Million-IRD dataset, additional visual results comparison between PromptIR (Potlapalli et al., 2023) and the proposed DyNet in all-in-one setting.
+Figure 8: Overview of the Transformer block used in our DyNet network.
 The Transformer block is composed of two sub-modules, the Multi Dconv head transposed attention module (MDTA) and the Gated Dconv feed-forward network (GDFN).
-A Transformer Block Details
+Appendix B Transformer Block Details
 As described in Section 3.1 of the main manuscript, here, we have discussed the transformer block utilized within our DyNet network architecture, detailing its sub-modules multi deconv head transposed attention (MDTA) and gated deconv feed-forward network (GDFN).
@@
 The activated feature map is then merged with the output from the alternative pathway via an element-wise multiplication.
-B Visual Results Comparison
-We show additional visual results comparison between PromptIR [36] and our DyNet-L under from all-in-one setting.
-Figure S2: Comparative analysis of image denoising by all-in-one methods on the BSD68 dataset [30]. Our DyNet-L reduces noise, producing more sharp and clear image compared to the PromptIR [36].
-Figure S3: Comparative analysis of image deraining by all-in-one methods on the Rain100L dataset [16].
-Our DyNet-L successfully eliminates rain streaks, producing clear, rain-free images as compared to the recent PromptIR [36].
-Figure S4: Comparative analysis of image dehazing by all-in-one methods on the SOTS dataset [25]. Our approach reduces haze, producing more clear image compared to the PromptIR [36].
-C Breakdown of images in Our Million-IRD Dataset
-We combine the existing high-quality, high-resolution natural image datasets such as LSDIR [27], DIV2K [2], Flickr2K [35], and NTIRE [4].
+Appendix C Visual Results Comparison
+We show additional visual results comparison between PromptIR (Potlapalli et al., 2023) and our DyNet under all-in-one setting.
+Figure 9: Comparative analysis of image denoising by all-in-one methods on the BSD68 dataset (Martin et al., 2001) and Urban100 (Huang et al., 2015).
+Our DyNet reduces noise, producing more sharper and clearer image compared to the PromptIR (Potlapalli et al., 2023).
+Figure 10: Comparative analysis of image deraining by all-in-one methods on the Rain100L dataset (Fan et al., 2019).
+Our DyNet-L successfully eliminates rain streaks, producing clear, rain-free images as compared to the recent PromptIR Potlapalli et al.
+(2023).
+Figure 11: Comparative analysis of image dehazing by all-in-one methods on the SOTS dataset (Li et al., 2018).
+Our approach reduces haze, producing more clear image compared to the PromptIR (Potlapalli et al., 2023).
+Appendix D Breakdown of images in Our Million-IRD Dataset
+We combine the existing high-quality, high-resolution natural image datasets such as LSDIR (Li et al., 2023a), DIV2K (Agustsson & Timofte, 2017), Flickr2K (Online, ), and NTIRE (Ancuti et al., 2021).
 Collectively these datasets have 90K images having spatial size ranging between $<1024^{2},4096^{2}>$.
-Their breakdown is given in Table S1.
+Their breakdown is given in Table 7.
 We also show the additional sample images from our Million-IRD dataset in Fig.
-S5 and Fig.
-S6.
-Table S1: Overview of the number of images curated from each database for our Million-IRD dataset.
-Dataset NTIRE [4] DIV2K [2] Flikr2K [35] LSDIR [27] Laion-HR [42]
+12 and Fig.
+13.
+Table 7: Overview of the number of images curated from each database for our Million-IRD dataset.
+Dataset NTIRE DIV2K Flikr2K LSDIR Laion-HR
+ (Ancuti et al., 2021)  (Agustsson & Timofte, 2017)  (Online, )  (Li et al., 2023a)  (Schuhmann et al., 2022)
 #Images 56 2,000 2,000 84,991 2M
-Figure S5: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs.
+Figure 12: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs.
 This includes a variety of textures, scenes from nature, sports activities, images taken during the day and at night, intricate textures, wildlife, shots captured from both close and distant perspectives, forest scenes, pictures of monuments, etc.
-Figure S6: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs.
+Figure 13: Sample images from our Million-IRD dataset, which features a diverse collection of high-quality, high-resolution photographs.
 This includes a variety of textures, scenes from nature, sports activities, images taken during the day and at night, intricate textures, wildlife, shots captured from both close and distant perspectives, forest scenes, pictures of monuments, etc.